OpenStudy (anonymous):

solve -sin(x)+cos(x)=0

OpenStudy (kc_kennylau):

Add sin(x) to both sides

OpenStudy (kc_kennylau):

What's the range of x?

OpenStudy (anonymous):

That just changes the format, the range of X is (0,2pi)

OpenStudy (kc_kennylau):

cos(x)=sin(x)

OpenStudy (kc_kennylau):

Change sin(x) to cos(pi/2-x)

OpenStudy (anonymous):

Why would you do that?

OpenStudy (kc_kennylau):

Because sin(x)=cos(pi/2-x)

OpenStudy (anonymous):

oh, so you would change sin(x) then solve it?

OpenStudy (kc_kennylau):



OpenStudy (kc_kennylau):

Yep

OpenStudy (anonymous):

Let me try that quick

OpenStudy (kc_kennylau):

Mind showing us your steps?

OpenStudy (anonymous):

well, now that I am thinking, I am not sure how you would solve it...

OpenStudy (kc_kennylau):

There are actually 3 methods to solve this

OpenStudy (kc_kennylau):

Method 1: cosx-sinx=0 cosx=sinx cosx=cos(pi/2-x) x=pi/2-x OR x=2pi+pi/2-x 2x=pi/2 OR 2x=5pi/2 x=pi/4 OR x=5pi/4 Method 2: cosx-sinx=0 cosx=sinx cos^2x=sin^2x cos^2x-sin^2x=0 cos(2x)=0 2x=pi/2 OR 2x=3pi/2 OR 2x=5pi/2 x=pi/4 OR x=3pi/4(rej) OR x=5pi/4 Method 3: cosx-sinx=0 cos^2x-2sinxcosx+sin^2x=0 -2sinxcosx=-1 sin(2x)=1 2x=pi/2 OR 2x=5pi/2 x=pi/4 OR 2x=5pi/4

OpenStudy (kc_kennylau):

Method 4 would be a modification of method 1 using sine instead of cosine

OpenStudy (ranga):

Or sin(x) = cos(x) tan(x) = 1 x = arctan(1)

OpenStudy (anonymous):

For what my problem is, method one would be best

OpenStudy (anonymous):

Thanks for your help guys, much appreciated

OpenStudy (kc_kennylau):

no problem